The motion of a particle is described by the equation u = at. The distance travelled by the particle in the first 4 second.

Grade:11

2 Answers

Kurma Anudeep
11 years ago

as u=at

replace u by v

v=at=10*4=40m/s

t=4s

d=s*t=40*4=160m

ravi sharma
6 years ago
At t=0, u=0
Now, use s=ut+1/2at^2
=> s= 0+1/2a(16)
        =8a
Done..............................................